calculate how many moles of product form when 1.75 mol of the reactant completely reacts 2S(s)+3O2 yields 2SO3(g)


1
Expert's answer
2022-04-05T00:49:03-0400

Mole ratio= 1:1

Mole produced= 1.75mol

Moles= mass/rfm

Moles of product= 1.75×1= 1.75